OpenStudy (anonymous):

What is the solution to this equation: 3x - 5 = 2x + 8 + x a. Identify (Infinity many solutions) b. No solution c. x = 0 d. x = 2

alones (alones):

simplify both sides of the equation.3x−5=2x+8+x 3x−5=3x+8 than subtract 3x from both sides. 3x−5−3x=3x+8−3x −5=8 and add 5 to both sides. −5+5=8+5 0=13 So i guess no solution :3